OpenStudy (anonymous):

Subtract. X^2+5x+6 / -(x^2+2x+1)

OpenStudy (anonymous):

i think its x+5/ -(x+1)

OpenStudy (ash2326):

@danni25 Is it subtraction or division?

OpenStudy (anonymous):

subtraction

OpenStudy (ash2326):

You mean this \[(x^2+5x+6)-(x^2+2x+1)\]?

OpenStudy (anonymous):

No its written in the book as x^2+5x+6 over -(x^2+2x+1)

OpenStudy (anonymous):

so if it gets reduced, its what i said before^^

OpenStudy (ash2326):

then it's just reduction, you need to factor the numerator and denominator and cancel the common terms @danni25

OpenStudy (anonymous):

anybody think they can help me on my question? c:

OpenStudy (ash2326):

We have this \[\large \frac{(x^2+5x+6)}{-(x^2+2x+1)}\] Let's factor the numerator first \[x^2+5x+6\] A standard quadratic equation is of the form \[ax^2+bx+c\] comparing the two we get a=1 b=6 c=5 now we need to find the factors of \(a\times c=5\times 1=5\) such that their sum is b (i.e 6) 1 and 5 satisfy this so we can write \[x^2+6x+5=x^2+5x+x+5\] Now it's easy to factor \[x(x+5)+1(x+5)\] Taking \(x+5\) common, we get \[(x+5)(x+1)\] Do you get this?

OpenStudy (anonymous):

yes thank you both

OpenStudy (ash2326):

No, it's not over yet. !!

OpenStudy (ash2326):

Can you do the rest?

OpenStudy (anonymous):

yes I just needed to know the steps to solving the beginning

OpenStudy (ash2326):

@danni25 Good:D